OpenStudy (anonymous):

An office clerk would require 48 hours to type a certain management report. A second clerk, who takes 3 minutes more to each page would require 60 hours. How many pages were there in the report and how long per page did it take each of them to type each page? Solution please too

OpenStudy (dumbcow):

60 - 48 = 12 It took 12 hours longer 12*60 = 720 It took 720 minutes longer 720/3 = 240 Therefore 240 pages 240/60 = 4 so 4 pages per hour OR 1 page every 15 minutes for the slower clerk 15-3 = 12 1 pg every 12 minutes for faster clerk

OpenStudy (anonymous):

if the difference between the two is 12 hours or 720 minutes. divided by 3 minutes for each pages is a total of 240 pages. then if it took 48 hours for the first one to do 240 pages that's 5 pages an hour or 20 minutes a page. and the second one is only 3 minutes more, so 23 minutes each page.